OpenStudy (anonymous):

How do you guys get the equations to look right when you are asking a question?

OpenStudy (anonymous):

i don't... maybe open another temporary window as a post-response, use the equation editor, and then cut-paste the \ [ ... \ ] source after you've prettied it up?

OpenStudy (anonymous):

thanks I will have to try that. It's kind of hard to ask for help if you can't write the equations down properly.

OpenStudy (saifoo.khan):

Try writing it in equation editor itself. Or use the draw tool.

OpenStudy (anonymous):

Openstudy uses something called MathJax in order to show the equations. MathJax is itself based on something known as LaTeX (pronounced Lay-tech) and it has it's own special terminology and symbology. For instance, in order to write, say \[\frac{u^3}{\sin(v)+2}\] all I need to write is "\frac{u^3}{\sin(v)+2}" and then enclose it in the \ [ and \ ] brackets mentioned above by Broken Fixer. Most math textbooks and papers are written using LaTeX, so it is something very familiar to mathematicians. Here is a link with a large number of notations used in LaTeX to write equations http://refcards.com/docs/silvermanj/tex/tex-refcard-letter.pdf Hope this is of help!

OpenStudy (saifoo.khan):